find the values of the trigonometric functions of theta from the information given. Sec theta =9, sin theta 《 0.a)sin theta_____ b)cos theta____c)csc theta__ d)sec theta__ e)cot theta__
Answer by stanbon(75887)

find the values of the trigonometric functions of theta from the information given. Sec theta =9, sin theta 《 0.
-----------------------
If sec is positive and sin is negative, theta is in QIV
where x is positive and y is negative.
-------------------------------
Since sec(t) = 9/1 = r/x, r = 9 and x = 1
Then y = -sqrt[9^2-1] = -sqrt(80) = -4sqrt(5)
--------------------------------------------------
a)sin theta = y/r = -4sqrt(5)/9
b)cos theta = 1/sec = 1/9
c)csc theta = y/x = -4sqrt(5)
d)sec theta = 9
e)cot theta = x/y = -1/4sqrt(5)
================================
Cheers,
Stan H.
================================